### v2

# Picocavity-enhanced Raman Spectroscopy of Physisorbed H2 and D2 Molecules

Abstract
We report on tip-enhanced Raman scattering (TERS) of H2 and D2 molecules physisorbed within a plasmonic picocavity at 10 K. The intense Raman peaks resulting from the rotational and vibrational transitions are observed at sub-nanometer gap distances of the junction formed by an Ag tip and Ag(111) surface, where a picocavity-enhanced field plays a crucial role. A significant red-shift of the H–H stretch frequency is found as the gap distance decreases, while the D–D stretch frequency is unaffected. Density functional theory, path-integral molecular dynamics, and quantum anharmonic vibrational energy calculations suggest that this unexpected isotope effect is explained by a different molecular density between H2 and D2 on the surface.
The adsorption of hydrogen molecules on solid surfaces is the first step in permeation for fuel storage [1], catalytic hydrogenation [2], hydrogen embrittlement [3], and nuclear-spin isomer conversion [4]. To understand and control the elementary processes in such physical and chemical phenomena, hydrogen adsorption has been intensively studied in surface science [5, 6, 7]. Specifically, hydrogen, the lightest molecule, exhibits unique quantum effects that differ from those of other molecules when it is physisorbed on a surface through van der Waals (vdW) interactions [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. However, only a few methods are available for characterizing such physisorption systems as the weak adsorption energy ( meV) necessitates measurements at cryogenic temperatures. Low-temperature scanning tunneling microscopy (LT-STM) has been employed as a local characterization technique to investigate the dynamics of hydrogen molecules weakly adsorbed on surfaces [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Several studies have observed characteristic features in conductance spectra, attributed to the rotational transition of molecular hydrogen in the STM junction [21, 22, 23, 24]. However, the interpretation remains controversial; the features within the corresponding energy range [30–60 meV (240–480 cm-1) for H2] could be attributed to configuration switching [18, 29, 30], to a molecule–substrate stretching mode [31], or to multiple phonon excitation [32]. Additionally, the stretching mode has never been observed in these studies.
Raman spectroscopy is a promising technique for characterizing hydrogen molecules on surfaces, as it allows the observation of both rotational and vibrational transitions of homonuclear diatomic molecules. Although the detection demands exceptional sensitivity due to the intrinsically small cross-section, single-molecule detection can be achieved by tip-enhanced and surface-enhanced Raman spectroscopy (TERS/SERS) [33, 34, 35, 36, 37]. Recent studies have established that picocavities, atomic-scale protrusions present in plasmonic nanojunctions, play a crucial role in achieving extreme confinement of the electromagnetic field to Ångström scales [38, 39, 37], resulting in a significant enhancement of Raman scattering [40, 41, 42] and ultrahigh spatial resolution [43, 44, 45, 46]. Yet, the exceptional sensitivity in TERS/SERS has been demonstrated primarily for organic molecules chemisorbed onto the substrate, where the enhancement benefits from chemical effects arising from the molecule–surface (and molecule–tip) interactions [47]. Therefore, applying TERS/SERS to physisorbed systems has remained challenging.
Here we demonstrate that LT-TERS can detect hydrogen molecules physisorbed onto an Ag(111) surface with an Ag tip [Fig. 1(a), inset], allowing the identification of their rotational and vibrational states. Furthermore, precise gap-distance control enables us to investigate not only the contribution of the picocavity field to the Raman scattering process but also the influence of tip proximity on the rotational/vibrational energies of the molecules through non-trivial nuclear quantum effects. Density functional theory (DFT), path-integral molecular dynamics (PIMD), and quantum vibrational energy calculations suggest that the tip–molecule–surface potential energy profiles are critically affected by the intermolecular and tip–molecular interactions, leading to unusual isotope effects on vibrational frequencies.
The LT-TERS experiments were performed in an ultra-high vacuum chamber [see Sect. I-A of the Supplemental Material (SM) [48]]. The 532-nm incident laser beam was linearly polarized along the tip axis and focused on the apex of a focused-ion-beam-sharpened Ag tip [40, 42] using an in-situ parabolic mirror, and the scattering light was detected by an ex-situ spectrometer. We conducted DFT calculations [49, 50] of H2/D2 on a Ag(111) (periodic) surface, employing the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional and screened pairwise vdW interactions [51] (see SM Sect. I-B [48]). The preferred adsorption sites (fcc/hcp hollow) and binding energies (30 meV) were in very good agreement with previous studies of isolated molecules on the surface [52, 53]. Unless specified otherwise, we considered a coverage of 0.69 monolayer (ML), motivated by the results of PIMD simulations at LT (see SM Sect. I-B [48]). At 0.69 ML coverage, the intermolecular attractive interaction increases the binding energy by 55 meV compared to the lowest coverage of 0.06 ML considered here (Fig. S4). Over the sample, we placed the Ag tip model shown in SM Sect. I-B [48] (see also Fig. S5), which was also previously discussed in Refs. [46, 54, 41].
H2 or D2 gas was introduced into the chamber where a clean Ag(111) surface was cooled at 10 K (see SM Sect. I-A [48]). Regardless of laser illumination, no static hydrogen molecule on the surface was imaged with LT-STM (Fig. S1). In contrast, clear peaks were observed in TERS spectra under these conditions. Figure 1(a) displays typical TERS spectra of H2 and D2 on Ag(111). For H2, intense peaks were observed at 351 and 4121 cm-1, along with a weak peak at 4443 cm-1. Similarly, for D2, intense peaks appeared at 169 and 2967 cm-1, with a weak peak at 3133 cm-1. The three peaks can be assigned to rotational/vibrotational modes of hydrogen. Note that broad or asymmetric peak shapes are artifacts in the wide wavenumber range spectra (see SM Sect. I-A [48]). The feature around 100 cm-1 common to both spectra originates from phonon modes of the Ag tip [41, 55], which is present even in the absence of hydrogen adsorption (Fig. S2). Figure 1(b) depicts the energy diagram of rotational/vibrational levels and allowed transitions under the Raman selection rule. The observed peaks, in order from lowest to highest, correspond to the lowest-energy rotational transition [; “rot” labeled in Fig. 1(a)], the lowest-energy vibrational transition [; “vib”], and their combination [; “rot+vib”], respectively. The three modes were also observed in high-resolution electron energy loss spectroscopy (HREELS) of hydrogen on a polycrystalline Ag film [8] and a Cu(001) surface [9] at approximately 10–12 K; however, better energy resolution (10 cm-1) in TERS than that of the HREELS (28 cm-1 [9]) makes the red-shifts from gas-phase hydrogen more discernible (Table S4). We should also note that only the stable nuclear spin isomer of each isotope, i.e., para-H2 and ortho-D2 [Fig. 1(b)], is detected in TERS (Fig. S2). On the surface, the stable isomer is dominant due to the ortho–para conversion [8, 4, 52] (see SM Sect. II-A [48]).
The rotational transition at 351 cm-1 for H2/Ag(111) [Fig. 1(a)] is slightly red-shifted from 354 cm-1 in the gas phase [56]. We calculate the H–H bonding distance to be 75.4 pm on Ag(111) (cf. 75.1 pm in the gas phase). Considering a rigid-rotor approximation, the rotational transition energy of para-H2 is calculated to be 353 cm-1 on the surface, compared to 356 cm-1 in the gas phase. This is in good agreement with the experiment. Notably, the H–H bonding distance of H2 on Ag(111) varied by only 0.06 pm at the closest tip position with respect to simulations without any tip. This suggests that the hydrogen molecules on the surface and in the STM junction behave as nearly free rotors.
In the TERS spectrum [Fig. 1(a)], the vibrational frequency of 4121 cm-1 exhibits a red-shifted relative to the gas-phase value of 4161 cm-1 [56]. The vibrational transition energy is calculated to be 4088 cm-1 on the surface (without the Ag tip), after scaling the calculations performed in the harmonic approximation (see SM Sect. I-C [48]; see also Fig. S9 for a simulated LT-TERS spectrum [46]). The slight discrepancy between the experimental and calculated frequencies are consistent with the typical spurious softening of the PBE functional. Compared to the calculated gas-phase value, the vibrational frequency shows a red-shift of 73 cm-1 upon the adsorption, which is larger than, but in reasonable agreement with the experiment.
The TERS intensity is strongly correlated with the localized surface plasmonic resonance (LSPR) of the tip–sample junction. In general, two types of enhancement mechanisms contribute to TERS [36, 47]: electromagnetic enhancement through LSPR and chemical enhancement originating from charge transfer between the target molecule and the tip/surface. In physisorption systems, the electromagnetic enhancement is expected to be dominant.
To elucidate the role of electromagnetic enhancement, it is crucial to investigate the gap-distance dependence of TERS signals. Figure 2(a) shows the waterfall plot of the gap-distance dependent TERS spectra recorded for D2/Ag(111). When the tip approaches the surface, the TERS intensities of the rotational and vibrational modes increase apparently exponentially [Fig. 2(b)], in agreement with the prediction based on the electric field of a light-induced dipole in the plasmonic tip [57]. A broad background with a maximum at 590 nm is attributed to electronic Raman scattering which represents the spectral response of the LSPR in the junction [58]. The gap-distance dependent intensities of both rotational and vibrational modes closely follow the plasmonic feature [Fig. 2(b)], indicating the dominant contribution of the gap-mode plasmon to the TERS enhancement. Additionally, we observed that changes in both the plasmon background and the molecular Raman peak occurred by an accidental modification of the tip-apex structure (Fig. S10), implying that the TERS enhancement is susceptible to the atomic-scale structure of the tip apex, which generates a plasmonic picocavity.
The gap-distance dependence of the TERS intensities is well reproduced by simulating the electronic field distribution in the Ag-tip–Ag-surface junction using the finite element method (FEM) (see SM [48] for details). To model the plasmonic picocavity, we assume an atomic-scale protrusion (0.5-nm radius of curvature) attached to a blunt tip body [40]. The black curve in Fig. 2(c) shows the field enhancement at the expected position of a D2 molecule located beneath the Ag tip, indicating a steep increase as the tip approaches. We confirmed the critical role of this sub-nanometric structure in TERS; as shown by the orange curve in Fig. 2(c), the tip without the atomic-scale protrusion exhibits a more moderate increase of the field enhancement as the tip approaches, which does not fit the experimental data [Fig. 2(b)]. The red curve in Fig. 2(c) shows the field enhancement at a possible position of the nearest neighboring molecule [Fig. 2(c), inset], indicating much lower enhancement. Therefore, TERS can achieve ‘nearly’ single-molecule sensitivity, even for physisorption systems. The contribution of LSPR to the TERS intensity was further confirmed by our excitation-wavelength dependent measurements (SM Sect. II-B [48]); a strong TERS peak was observed only when the scattering light wavelength matches with the resonance range of the LSPR (Fig. S3).
TERS reveals the sensitivity of the rotational/vibrational states to the interactions between a target molecule and its local environment. Figures 3(a)–3(d) show the gap-distance dependence of high-resolution TERS spectra focusing on the rotational/vibrational modes of H2/D2. Notably, only the vibrational frequency of H2 exhibits a significant red-shift as the gap distance decreases [Fig. 3(c)]; H2 shows a peak shift of 33 cm-1 with a 200 pm displacement of the tip, whereas is merely cm-1 for D2. Additionally, the peak width of H2 vibrational mode also increases as the gap distance decreases (see SM Sect. II-C [48] for details). In contrast, the peak positions of the rotational modes for both H2 [Fig. 3(a)] and D2 [3(b)] are almost unchanged, consistent with our calculations showing that the interatomic distance is independent of the tip height, as described above.
The isotope dependence of the red-shift of the vibrational band far exceeds the scale of typical isotope effects. To understand this notable behavior, we calculated the potential energy profile along the rigid vertical displacement of the center of mass of one molecule between the surface and the tip apex [Fig. 4(a)], while varying the tip–sample distance [see profiles (i)–(iv) at different in Fig. 4(b)]. At each position, we calculated the H2 and D2 harmonic intramolecular vibrational frequency (see SM Sect. II-C [48]). The calculations indicate a pronounced coupling of the stretch coordinate with the molecular translation along , evidenced by a steep variation of along this coordinate (Fig. S8). Furthermore, this variation changes depending on the tip–sample distance, leading us to define . As detailed in SM Sect. I-C [48], for each , we constructed a model 2D Schrödinger equation in the coupled coordinates and , parameterized by and . The solution of this model gives us the expected intramolecular stretch vibrational transition energy for each isotope, fully accounting for the anharmonic coupling between these coordinates. We concluded that if both H2 and D2 probe the same well of [Fig. 4(b)], the isotope effect on can account at most for a factor 2 between H2 and D2 (Tables S2 and S3), in disagreement with the experiment (Fig. 3).
We demonstrate that this discrepancy can be resolved by accounting for density differences between H2 and D2. At lower temperatures, the solids formed by H2 and D2 have different molar volumes due to nuclear quantum effects [zero-point energy (ZPE) swelling [59]; see also SM Sect. I-B [48]]. As a result, solid D2 is more than twice as dense as solid H2 at 4.2 K [60]. Because the interaction potential between the molecules and the surface exhibits almost no corrugation [52], we expect that H2 adsorbs on Ag(111) at 10 K more sparsely than D2. To explore a limiting low-density regime, we repeated the calculations discussed in the previous paragraph for a coverage of 0.06 ML of H2 [Fig. 4(c)]. As shown in Fig. 4(d), the reduced intermolecular interactions at the surface cause the minimum of along to lie closer to the tip instead of closer to the surface, as decreases [(i) to (iv)]. Consequently, the average position of the molecule center of mass [bullet points in Figs. 4(b) and (d)] also shifts towards the tip as decreases [dotted line in Fig. 4(d)]. Computing in this scenario with the 2D quantum model of and , we find a larger variation with than in the higher-density case (Table S2). As shown in Fig. 4(e), considering the higher-density scenario for D2 (the filled bullets) and the lower-density scenario for H2 (the empty bullets) can account for the large isotope effect in the red-shift of the stretch transition as the tip approaches. Figure 4(e) also shows the experimental values from the plots in Fig. 3 (the solid and dotted curves for H2 and D2, respectively; see also Fig. S11). Although the exact gap distance in STM experiment is hard to determine (see SM Sect. II-C [48]), the overlap of the experimental curves with the calculated plots in Fig. 4(e) shows good agreement. The remaining numerical discrepancies can be attributed to the limitation of the DFT level of theory (see SM Sect. I-B [48]), the exact tip apex geometry, and to the uncertainty in determining the exact local surface coverage. Nevertheless, our results highlight the crucial role of intermolecular interactions in accurately interpreting vibrational spectroscopy of hydrogen isotopes in nanojunctions, where nuclear quantum effects have a significant impact [13].
In summary, with LT-TERS, we observed the rotational/vibrational transitions of local H2 and D2 molecules through the interaction with the picocavity field. Our results show that LT-TERS is applicable to weakly adsorbed molecules, offering deeper insights into the local structures, reactions, and dynamics of adsorbates, such as chemical reactions at active sites [61] and surface diffusion (spillover) in catalytic processes [62], at the single-molecule level.
Acknowledgements.
We thank Adnan Hammud for the Ag-tip fabrication. We thank Jun Yoshinobu and Heiko Appel for helpful discussions. T.S. acknowledges the support of JST FOREST Grant No. JPMJFR221U. T.K. acknowledges the support of JST FOREST Grant No. JPMJFR201J and JSPS KAKENHI Grant No. JP19K24684.References
- Ockwig and Nenoff [2007] N. W. Ockwig and T. M. Nenoff, Membranes for hydrogen separation, Chem. Rev. 107, 4078 (2007).
- Zhang et al. [2019] L. Zhang, M. Zhou, A. Wang, and T. Zhang, Selective hydrogenation over supported metal catalysts: from nanoparticles to single atoms, Chem. Rev. 120, 683 (2019).
- Li et al. [2020] X. Li, X. Ma, J. Zhang, E. Akiyama, Y. Wang, and X. Song, Review of hydrogen embrittlement in metals: hydrogen diffusion, hydrogen characterization, hydrogen embrittlement mechanism and prevention, Acta Metallurgica Sinica 33, 759 (2020).
- Fukutani and Sugimoto [2013] K. Fukutani and T. Sugimoto, Physisorption and ortho–para conversion of molecular hydrogen on solid surfaces, Prog. Surf. Sci 88, 279 (2013).
- Christmann [1988] K. Christmann, Interaction of hydrogen with solid surfaces, Surf. Sci. Rep. 9, 1 (1988).
- Vidali et al. [1991] G. Vidali, G. Ihm, H.-Y. Kim, and M. W. Cole, Potentials of physical adsorption, Surf. Sci. Rep. 12, 135 (1991).
- Bruch et al. [2007] L. Bruch, R. Diehl, and J. Venables, Progress in the measurement and modeling of physisorbed layers, Rev. Mod. Phys. 79, 1381 (2007).
- Avouris et al. [1982] P. Avouris, D. Schmeisser, and J. Demuth, Observation of rotational excitations of H2 adsorbed on Ag surfaces, Phys. Rev. Lett. 48, 199 (1982).
- Andersson and Harris [1982] S. Andersson and J. Harris, Observation of rotational transitions for H2, D2, and HD adsorbed on Cu(100), Phys. Rev. Lett. 48, 545 (1982).
- Yu et al. [1983] C.-f. Yu, K. B. Whaley, C. S. Hogg, and S. J. Sibener, Selective adsorption resonances in the scattering of -H2, -H2, -D2 and -D2 from Ag(111), Phys. Rev. Lett. 51, 2210 (1983).
- Sakurai et al. [1988] M. Sakurai, T. Okano, and Y. Tuzi, Ortho-para conversion of -H2 physisorbed on Ag(111) near two-dimensional condensation conditions, Appl. Surf. Sci. 33, 245 (1988).
- Anger et al. [1989] G. Anger, A. Winkler, and K. Rendulic, Adsorption and desorption kinetics in the systems H2/Cu(111), H2/Cu(110) and H2/Cu(100), Surf. Sci. 220, 1 (1989).
- Williams et al. [2002] K. A. Williams, B. K. Pradhan, P. C. Eklund, M. K. Kostov, and M. W. Cole, Raman spectroscopic investigation of H2, HD, and D2 physisorption on ropes of single-walled, carbon nanotubes, Phys. Rev. Lett. 88, 165502 (2002).
- Fukutani et al. [2003] K. Fukutani, K. Yoshida, M. Wilde, W. Diño, M. Matsumoto, and T. Okano, Photostimulated desorption and ortho-para conversion of H2 on Ag surfaces, Phys. Rev. Lett. 90, 096103 (2003).
- Niki et al. [2008] K. Niki, T. Kawauchi, M. Matsumoto, K. Fukutani, and T. Okano, Mechanism of the ortho-para conversion of hydrogen on Ag surfaces, Phys. Rev. B 77, 201404(R) (2008).
- Sugimoto and Fukutani [2014] T. Sugimoto and K. Fukutani, Effects of rotational-symmetry breaking on physisorption of ortho-and para-H2 on Ag(111), Phys. Rev. Lett. 112, 146101 (2014).
- Sugimoto et al. [2017] T. Sugimoto, Y. Kunisada, and K. Fukutani, Inelastic electron tunneling mediated by a molecular quantum rotator, Phys. Rev. B 96, 241409(R) (2017).
- Gupta et al. [2005] J. Gupta, C. Lutz, A. Heinrich, and D. Eigler, Strongly coverage-dependent excitations of adsorbed molecular hydrogen, Phys. Rev. B 71, 115416 (2005).
- Temirov et al. [2008] R. Temirov, S. Soubatch, O. Neucheva, A. C. Lassise, and F. S. Tautz, A novel method achieving ultra-high geometrical resolution in scanning tunnelling microscopy, New J. Phys. 10, 053012 (2008).
- Lotze et al. [2012] C. Lotze, M. Corso, K. J. Franke, F. von Oppen, and J. I. Pascual, Driving a macroscopic oscillator with the stochastic motion of a hydrogen molecule, Science 338, 779 (2012).
- Li et al. [2013] S. Li, A. Yu, F. Toledo, Z. Han, H. Wang, H. He, R. Wu, and W. Ho, Rotational and vibrational excitations of a hydrogen molecule trapped within a nanocavity of tunable dimension, Phys. Rev. Lett. 111, 146102 (2013).
- Natterer et al. [2013] F. D. Natterer, F. Patthey, and H. Brune, Distinction of nuclear spin states with the scanning tunneling microscope, Phys. Rev. Lett. 111, 175303 (2013).
- Natterer et al. [2014] F. D. Natterer, F. Patthey, and H. Brune, Resonant-enhanced spectroscopy of molecular rotations with a scanning tunneling microscope, ACS Nano 8, 7099 (2014).
- Therrien et al. [2015] A. J. Therrien, A. Pronschinske, C. Murphy, E. A. Lewis, M. L. Liriano, M. D. Marcinkowski, and E. C. H. Sykes, Collective effects in physisorbed molecular hydrogen on Ni/Au(111), Phys. Rev. B 92, 161407(R) (2015).
- Liu et al. [2018] S. Liu, A. Shiotari, D. Baugh, M. Wolf, and T. Kumagai, Enhanced resolution imaging of ultrathin ZnO layers on Ag(111) by multiple hydrogen molecules in a scanning tunneling microscope junction, Phys. Rev. B 97, 195417 (2018).
- Merino et al. [2018] P. Merino, A. Rosławska, C. C. Leon, A. Grewal, C. Große, C. González, K. Kuhnke, and K. Kern, A single hydrogen molecule as an intensity chopper in an electrically driven plasmonic nanocavity, Nano Lett. 19, 235 (2018).
- Wang et al. [2022] L. Wang, Y. Xia, and W. Ho, Atomic-scale quantum sensing based on the ultrafast coherence of an H2 molecule in an STM cavity, Science 376, 401 (2022).
- Wang et al. [2023] L. Wang, D. Bai, Y. Xia, and W. Ho, Electrical manipulation of quantum coherence in a two-level molecular system, Phys. Rev. Lett. 130, 096201 (2023).
- Thijssen et al. [2006] W. Thijssen, D. Djukic, A. Otte, R. Bremmer, and J. van Ruitenbeek, Vibrationally induced two-level systems in single-molecule junctions, Phys. Rev. Lett. 97, 226806 (2006).
- Halbritter et al. [2008] A. Halbritter, P. Makk, S. Csonka, and G. Mihály, Huge negative differential conductance in Au–H2 molecular nanojunctions, Phys. Rev. B 77, 075402 (2008).
- Li et al. [2015] Y. Li, S. Kaneko, S. Fujii, and M. Kiguchi, Symmetry of single hydrogen molecular junction with Au, Ag, and Cu electrodes, J. Phys. Chem. C 119, 19143 (2015).
- Trouwborst et al. [2009] M. Trouwborst, E. Huisman, S. van der Molen, and B. van Wees, Bistable hysteresis and resistance switching in hydrogen-gold junctions, Phys. Rev. B 80, 081407(R) (2009).
- Haran [2010] G. Haran, Single-molecule Raman spectroscopy: a probe of surface dynamics and plasmonic fields, Acc. Chem. Res. 43, 1135 (2010).
- Ding et al. [2016] S.-Y. Ding, J. Yi, J.-F. Li, B. Ren, D.-Y. Wu, R. Panneerselvam, and Z.-Q. Tian, Nanostructure-based plasmon-enhanced Raman spectroscopy for surface analysis of materials, Nat. Rev. Mater. 1, 16021 (2016).
- Wang et al. [2020] X. Wang, S.-C. Huang, S. Hu, S. Yan, and B. Ren, Fundamental understanding and applications of plasmon-enhanced Raman spectroscopy, Nat. Rev. Phys. 2, 253 (2020).
- Itoh et al. [2023] T. I. Itoh, M. Procházka, Z.-C. Dong, W. Ji, Y. S. Yamamoto, Y. Zhang, and Y. Ozaki, Toward a new era of SERS and TERS at the nanometer scale: From fundamentals to innovative applications, Chem. Rev. 123, 1552 (2023).
- Höppener et al. [2024] C. Höppener, J. Aizpurua, H. Chen, S. Gräfe, A. Jorio, S. Kupfer, Z. Zhang, and V. Deckert, Tip-enhanced Raman scattering, Nat. Rev. Methods Primers 4, 47 (2024).
- Benz et al. [2016] F. Benz, M. K. Schmidt, A. Dreismann, R. Chikkaraddy, Y. Zhang, A. Demetriadou, C. Carnegie, H. Ohadi, B. De Nijs, R. Esteban, J. Aizpurua, and J. J. Baumberg, Single-molecule optomechanics in “picocavities”, Science 354, 726 (2016).
- Urbieta et al. [2018] M. Urbieta, M. Barbry, Y. Zhang, P. Koval, D. Sánchez-Portal, N. Zabala, and J. Aizpurua, Atomic-scale lightning rod effect in plasmonic picocavities: A classical view to a quantum effect, ACS Nano 12, 585 (2018).
- Liu et al. [2019] S. Liu, M. Müller, Y. Sun, I. Hamada, A. Hammud, M. Wolf, and T. Kumagai, Resolving the correlation between tip-enhanced resonance raman scattering and local electronic states with 1 nm resolution, Nano Lett. 19, 5725 (2019).
- Liu et al. [2023] S. Liu, F. P. Bonafe, H. Appel, A. Rubio, M. Wolf, and T. Kumagai, Inelastic light scattering in the vicinity of a single-atom quantum point contact in a plasmonic picocavity, ACS Nano 17, 10172 (2023).
- Park et al. [2024] Y. Park, I. Hamada, A. Hammud, T. Kumagai, M. Wolf, and A. Shiotari, Atomic-precision control of plasmon-induced single-molecule switching in a metal–semiconductor nanojunction, Nat. Commun. 15, 6709 (2024).
- Zhang et al. [2013] R. Zhang, Y. Zhang, Z. Dong, S. Jiang, C. Zhang, L. Chen, L. Zhang, Y. Liao, J. Aizpurua, Y. Luo, J. L. Yang, and J. G. Hou, Chemical mapping of a single molecule by plasmon-enhanced Raman scattering, Nature 498, 82 (2013).
- Lee et al. [2019] J. Lee, K. T. Crampton, N. Tallarida, and V. A. Apkarian, Visualizing vibrational normal modes of a single molecule with atomically confined light, Nature 568, 78 (2019).
- Xu et al. [2021] J. Xu, X. Zhu, S. Tan, Y. Zhang, B. Li, Y. Tian, H. Shan, X. Cui, A. Zhao, Z. Dong, J. Yang, Y. Luo, B. Wang, and J. G. Hou, Determining structural and chemical heterogeneities of surface species at the single-bond limit, Science 371, 818 (2021).
- Litman et al. [2023] Y. Litman, F. P. Bonafé, A. Akkoush, H. Appel, and M. Rossi, First-principles simulations of tip enhanced Raman scattering reveal active role of substrate on high-resolution images, J. Phys. Chem. Lett. 14, 6850 (2023).
- Chen and Jensen [2023] R. Chen and L. Jensen, Interpreting chemical enhancements of surface-enhanced Raman scattering, Chem. Phys. Rev. 4 (2023).
- [48] See Supplemental Material for detailed experimental and theoretical methods, additional figures, and tables, which includes Refs. [63, 40, 49, 50, 51, 64, 46, 41, 52, 65, 17, 66, 67, 54, 68, 57, 69, 4, 56, 8, 11, 14, 15, 9, 70, 71, 18].
- Blum et al. [2009] V. Blum, R. Gehrke, F. Hanke, P. Havu, V. Havu, X. Ren, K. Reuter, and M. Scheffler, Ab initio molecular simulations with numeric atom-centered orbitals, Comput. Phys. Commun. 180, 2175 (2009).
- Gavini et al. [2023] V. Gavini, S. Baroni, V. Blum, D. R. Bowler, A. Buccheri, J. R. Chelikowsky, S. Das, W. Dawson, P. Delugas, M. Dogan, C. Draxl, G. Galli, L. Genovese, P. Giannozzi, M. Giantomassi, X. Gonze, M. Govoni, F. Gygi, A. Gulans, J. M. Herbert, S. Kokott, T. D. Kühne, K.-H. Liou, T. Miyazaki, P. Motamarri, A. Nakata, J. E. Pask, C. Plessl, L. E. Ratcliff, R. M. Richard, M. Rossi, R. Schade, M. Scheffler, O. Schütt, P. Suryanarayana, M. Torrent, L. Truflandier, T. L. Windus, Q. Xu, V. W.-Z. Yu, and D. Perez, Roadmap on electronic structure codes in the exascale era, Model. Simul. Mat. Sci. Eng. 31, 063301 (2023).
- Ruiz et al. [2012] V. G. Ruiz, W. Liu, E. Zojer, M. Scheffler, and A. Tkatchenko, Density-functional theory with screened van der Waals interactions for the modeling of hybrid inorganic-organic systems, Phys. Rev. Lett. 108, 146103 (2012).
- Kunisada and Kasai [2015] Y. Kunisada and H. Kasai, Hindered rotational physisorption states of H2 on Ag(111) surfaces, Phys. Chem. Chem. Phys. 17, 19625 (2015).
- Smeets and Kroes [2021] E. W. Smeets and G.-J. Kroes, Designing new SRP density functionals including non-local vdW-DF2 correlation for H2 + Cu(111) and their transferability to H2 + Ag(111), Au(111) and Pt(111), Phys. Chem. Chem. Phys. 23, 7875 (2021).
- Cirera et al. [2022] B. Cirera, Y. Litman, C. Lin, A. Akkoush, A. Hammud, M. Wolf, M. Rossi, and T. Kumagai, Charge transfer-mediated dramatic enhancement of Raman scattering upon molecular point contact formation, Nano Lett. 22, 2170 (2022).
- Cui et al. [2023] J. Cui, Y. Zhang, and Z.-C. Dong, Atomistic polarization model for Raman scattering simulations of large metal tips with atomic-scale protrusions at the tip apex, Nanophotonics 12, 4031 (2023).
- Veirs and Rosenblatt [1987] D. K. Veirs and G. M. Rosenblatt, Raman line positions in molecular hydrogen: H2, HD, HT, D2, DT, and T2, J. Mol. Spectrosc. 121, 401 (1987).
- Pettinger et al. [2009] B. Pettinger, K. F. Domke, D. Zhang, G. Picardi, and R. Schuster, Tip-enhanced Raman scattering: influence of the tip-surface geometry on optical resonance and enhancement, Surf. Sci. 603, 1335 (2009).
- Kamimura et al. [2022] R. Kamimura, T. Kondo, K. Motobayashi, and K. Ikeda, Surface-enhanced electronic Raman scattering at various metal surfaces, Phys. Status Solidi B 259, 2100589 (2022).
- Markland et al. [2008] T. E. Markland, S. Habershon, and D. E. Manolopoulos, Quantum diffusion of hydrogen and muonium atoms in liquid water and hexagonal ice, J. Chem. Phys. 128, 194506 (2008).
- Megaw and Simon [1936] H. D. Megaw and F. Simon, Density and compressibility of solid hydrogen and deuterium at 4.2 K, Nature 138, 244 (1936).
- Cai et al. [2023] Z.-F. Cai, N. Kumar, and R. Zenobi, Probing on-surface chemistry at the nanoscale using tip-enhanced Raman spectroscopy, CCS Chem. 5, 55 (2023).
- Yang et al. [2024] J.-L. Yang, Y.-W. Zhou, M.-Y. Jia, Q. Liu, and Z.-F. Cai, Insights into heterogeneous catalysis by tip-enhanced Raman spectroscopy, ChemCatChem 16, e202400370 (2024).
- Summers [1969] R. L. Summers, Empirical observations on the sensitivity of hot cathode ionization type vacuum gages, NASA Technical Report NASA-TN-D-5285 (1969).
- Hermann and Tkatchenko [2020] J. Hermann and A. Tkatchenko, Density functional model for van der Waals interactions: Unifying many-body atomic approaches with nonlocal functionals, Phys. Rev. Lett. 124, 146401 (2020).
- Kapil et al. [2019] V. Kapil, M. Rossi, O. Marsalek, R. Petraglia, Y. Litman, T. Spura, B. Cheng, A. Cuzzocrea, R. H. Meiß ner, D. M. Wilkins, B. A. Helfrecht, P. Juda, S. P. Bienvenue, W. Fang, J. Kessler, I. Poltavsky, S. Vandenbrande, J. Wieme, C. Corminboeuf, T. D. Kühne, D. E. Manolopoulos, T. E. Markland, J. O. Richardson, A. Tkatchenko, G. A. Tribello, V. V. Speybroeck, and M. Ceriotti, i-PI 2.0: A universal force engine for advanced molecular simulations, Comput. Phys. Commun. 236, 214 (2019).
- Schwartz [1985] C. Schwartz, High-accuracy approximation techniques for analytic functions, J. Math. Phys. 26, 411 (1985).
- Colbert and Miller [1992] D. T. Colbert and W. H. Miller, A novel discrete variable representation for quantum mechanical reactive scattering via the -matrix Kohn method, J. Chem. Phys. 96, 1982 (1992).
- Jiang et al. [2016] Y. Jiang, S. Pillai, and M. A. Green, Realistic silver optical constants for plasmonics, Sci. Rep. 6, 30605 (2016).
- Wolniewicz [1983] L. Wolniewicz, The state vibration-rotational energies of the H2, HD, and D2 molecules, J. Chem. Phys. 78, 6173 (1983).
- Teal and MacWood [1935] G. K. Teal and G. E. MacWood, The Raman spectra of the isotopic molecules H2, HD, and D2, J. Chem. Phys. 3, 760 (1935).
- Persson and Ryberg [1989] B. Persson and R. Ryberg, Vibrational line shapes of low-frequency adsorbate modes: CO on Pt(111), Phys. Rev. B 40, 10273 (1989).
Supplemental Material: Picocavity-enhanced Raman Spectroscopy of Physisorbed H2 and D2 Molecules
I I. Methods
I.1 A. STM-TERS Experiments
The experiments were performed in an ultrahigh vacuum (UHV) chamber equipped with an STM-based TERS machine (modified UNISOKU USM-1400) at a sample temperature of 10 K. A single-crystalline Ag(111) surface (MaTeck) was cleaned by Ar-sputtering-and-annealing cycles in the chamber. H2 (99.999%, Westfalen AG) or D2 (99.7%, Air Liquide) gas was introduced into the STM unit cooled at 10 K via a variable leak valve. The gas exposure (Langmuir; L) is estimated by , where [ Torr], [s], and denote the displayed pressure value of the gauge (corrected by N2), the exposure duration, and the correction factor for the gauge, respectively. A Bayard-Alpert ionization gauge is located outside the cryogen-cooled STM unit in the chamber and we used for H2 and for D2 [63]. We used a chemically-etched Ag tip that was further sharpened by focused-ion-beam milling as described elsewhere [40]. The bias voltage was applied to the sample while the tip was grounded. The STM images were obtained in the constant current mode at = 50 pA. The cleanness of the Ag(111) surface and tip was confirmed by STM images [Fig. S1(a)]. We also confirmed that a TERS spectrum of the Ag–Ag junction before hydrogen dosing had no molecule-derived peaks (Fig. S2, at 0 L).
In the STM junction, -polarised visible laser (532-nm solid-state or 633-nm HeNe) was directed through a fused silica window and an Ag-coated parabolic mirror (numerical aperture of 0.6). The incident laser power is 5.9 mW at = 532 nm and 9.7 mW at 633 nm. The parabolic mirror is mounted on a five-axis piezoelectric-motor stage in the LT-STM unit and the laser is focused at the tip apex by adjusting the position and angle of the parabolic mirror. The scattered light collected by the parabolic mirror was directed to a spectrometer (AndorShamrock 303i) outside the UHV chamber via a beam splitter and a longpass filter, each selected according to the excitation wavelength. For the TERS and STM-induced luminescence (STML) spectra, we used the gratings 150 (for Fig. S3), 600 (Figs. 1, 2, and S2), and 1200 line/mm (Fig. 3 and S10). A lower-density grating allows the acquisition of spectra over a wider wavelength range, as shown in Fig. S3 (see Sect. II-B), but resulting in broad and asymmetric Raman peaks. Therefore, for the high-frequency-resolution measurements shown in Fig. 3, the densest grating was used.
I.2 B. DFT Calculations and Molecular Dynamics
DFT calculations were performed with the FHI-aims program package [49, 50]. We used intermediate defaults for basis sets, radial grids and cutoff potential parameters and light defaults for the Hartree multipole expansion and angular grids. This combination proved efficient and reliable against benchmarks to tight settings. We simulated an orthogonal Ag(111) surface cell oriented perpendicular to the axis, containing four layers. The cell parameters used were Å, Å and Å. We employed a -point grid and a dipole correction was employed to maximally decouple periodic images along the direction. We used the PBE exchange-correlation functional and included vdW interactions through the pairwise screened TS-vdWsurf method of Ref. [51] in all simulations reported in the main text. Figure S4 shows a comparison of binding energies with the many-body MBD-NL dispersion correction of Ref. [64], indicating a weaker binding energy by only 5 meV and a binding distance 0.1 Å closer to the surface. Structures were optimized by allowing only hydrogen molecules and the top layer of Ag atoms of the surface to relax. All the other atoms were kept fixed at their bulk (or tip) positions. When we performed simulations including a model tip structure, this was modeled by an Ag19 cluster at varying distances from the surface. The geometry of the tip is shown in Fig. S5. Although this tip size would not be sufficient to capture a proper plasmonic response of the tip, the tip structure can capture the local interactions with molecules close to the tip apex [46, 54, 41].
We inspected binding sites and binding energies of a single adsorbed H2 molecule, finding all values in good agreement with previous studies [52]. The binding energy was defined as the difference between the total energy when the molecule was 10 Å away from the surface and the minimum total energy of the adsorbed molecule. The blue curve in Fig. S4 shows the binding curve of a single H2 molecule adsorbed on the fcc-hollow site of the Ag(111) surface, indicating an equilibrium H2–Ag(111) bonding length of 3.09 Å and a binding energy of 32 meV. The model we used in most simulations contained 11 molecules in the unit cell (see rationale below). For this molecular concentration, we evaluated the intermolecular interaction by comparing the binding curves (orange curve in Fig. S4) with that of the single-molecule case (blue); the molecule–molecule interaction in our models amounts to 55 meV/molecule.
Performing vibrational analysis on the optimized single molecule adsorbed on the surface showed that negative frequencies related to rotational and translational modes of the molecule were always present. This led us to consider higher coverages to model local stable clustering of the molecules on the surface. We considered a full monolayer (1 ML), where all fcc-hollow sites of the unit cell were occupied by H2. With this setup we performed Born-Oppenheimer molecular dynamics (BOMD) and PIMD at LT ( = 20 K) with the i-PI code [65] connected to FHI-aims. PIMD simulations used 12 beads in conjunction with a normal-mode colored-noise thermostat with 8 auxiliary variables and . This number of beads is not fully converged at this temperature even with the colored-noise quantum thermostats, but the purpose of these simulations was exploratory.
While the full monolayer was stable during more than 10 ps on the classical trajectories, in the quantum trajectory 5 molecules readily desorbed (within the first picosecond) from the substrate. For the quantum case, we validated this observation by simulating trajectories with a parameterized Lennard-Jones potential representing the Ag(111) surface. We could calculate the radius of gyration of H atoms from the PIMD simulations, belonging to the 3 molecules closest to the tip at any point in the simulation. For the closest tip distance considered ( = 7.4 Å), these were 0.33 Å along and and 0.30 Å along ( = 20 K), indicating a small anisotropy. Converting these numbers to an effective thermal de Broglie wavelength and applying a inverse-square-root scaling for the temperature and mass, we would obtain Å for H2 and Å for D2 at 10 K. We note that, depending on the density of H2 molecules on the surface, for H2 is about the length of the inter-particle spacing at 10 K and signatures of bosonic exchange could begin to be observed. These features are disregarded in the current simulations. Further, we note that even though these simulations provide a clear indication that quantum mechanical effects affect the density of the molecular coverage of the surface, a complete determination would require much longer simulations, which would take years to complete or require the development of a tailored machine-learning potential that surpasses the scope of this work.
Figure S6 shows the energy profiles experienced by an H2 molecule when varying the distance of its center of mass with respect to the Ag(111) surface, at different tip-surface distances, and at 0.69 ML coverage. We only varied the position of the molecule closest to the tip apex, while the others were kept at equilibrium. The profile shapes are similar to an asymmetric double-well potential at closer tip–surface distances, while they assume a shape closer to a Morse potential at larger distances, which asymptotically approaches the profile of an H2 molecule on Ag(111) without the Ag tip (yellow curve in Fig. S6). The asymmetry of the double-well is strongly impacted by the presence of other molecules on the surface (molecule–molecule interactions). Figure S7(a) shows a comparison with the limiting case of a single molecule in the unit cell (0.06 ML coverage) at a tip–surface distance of 9.4 Å. In that case (orange curve), if the molecule can overcome the barrier, it will have a higher probability of being found closer to the tip, which was predicted in a Au-tip–single-H2–Au-surface junction [17]. As we discuss in the main text, this effect is essential to explain the isotope-dependent red-shift observed experimentally.
Figure S7(b) shows a comparison of the 0.69 ML coverage with the tip fixed at Å and different vdW dispersion corrections, added to the PBE functional. It is clear that details of the vdW dispersion impact the shape of the potential. In this case, the many-body vdW dispersion decreases the distance between the two minima of the asymmetric double-well
I.3 C. Model 2D Schrödinger Equation
[…]
where is the potential energy of a molecule with its center of mass fixed at a distance above the substrate surface and with all other coordinates kept at their equilibrium values, as described at the end of Sec. I.2. The vibrational motion is parametrized by the reduced mass and angular frequency . We set the mass of a hydrogen atom to and use .
The vibrational frequencies are computed from numerical (finite-difference) second derivatives of the DFT energies with respect to the vibrational coordinate,
[…]
We computed the numerically exact solutions to the 2D Schrödinger equation corresponding to Eq. (S1) using the sinc-function discrete variable representation (DVR) [66, 67]. The equation was solved for high coverage (0.69 ML) with and . Given the large separation in the timescales of translational and vibrational motion, we also considered the vibrationally adiabatic approximation, whereby one fixes the translational coordinate in Eq. (S1) and solves the corresponding one-dimensional (1D) Schrödinger equation in , yielding the analytical harmonic vibrational energies
These are then used to construct the vibrationally adiabatic translational potential,
for which we solve the 1D Schrödinger equation numerically using DVR. The energy difference between the ground state solutions for and is an excellent approximation to the vibrational excitation energy of the 2D model (see Table S1). Given the accuracy of the vibrationally adiabatic approximation, we adopt it in all subsequent calculations.
2D | ||||
1D |
[…]
11.4 | ||||||
9.4 | ||||||
7.9 | ||||||
7.4 |
11.4 | ||||||
9.4 | ||||||
7.9 | ||||||
7.4 |
I.4 D. TERS Spectrum Simulation
We conducted TERS simulations according to the methodology proposed in Refs. [46]. We used Tip A discussed in Ref. [54] and calculated the anharmonic Raman spectrum based on an ab initio molecular dynamics (AIMD) trajectory (5 ps) of a molecular layer adsorbed on the surface and a tip apex located 4 Å above the molecular layer at = 20 K. We calculated the TERS intensity as , where was obtained by a local polarizability calculation at each geometry in the time-series given by the molecular dynamics. For this part of the calculation, we only considered the free-standing molecular layer (no surface) at the positions adopted in the AIMD trajectory including the surface. We note that the method of Ref. [46] is unable to capture the plasmonic enhancement in its totality. We do not observe any noteworthy features in the calculated spectrum and it overall agrees with the experimental spectrum. The red-shift of the rotational band in the simulation is explained by the absence of nuclear spin in the simulations, making calculations reproduce only classical rotations that are red-shifted compared to the quantized transition of para-H2. This simulation is of exploratory character, as achieving statistical convergence through the use of parametrized machine-learned potentials along with accurate plasmonic enhancements is beyond the scope of this paper.
I.5 E. FEM Simulations
To evaluate the electric field in the STM junction, we performed FEM simulations using COMSOL Multiphysics (version 6.2 with the Wave Optics Module). We used the three-dimentional model with an Ag tip and Ag sample plate described in Supporting Information of Ref. [40]. To reproduce a plasmonic picocavity in the STM junction, at the apex of the Ag tip (a length of 300 nm, an effective radius of 30 nm), an Ag half sphere with a radius of 0.5 nm is attached as an atomic-scale protrusion. For comparison, we also constructed another Ag tip model with an effective radius of 30 nm without the atomic-scale protrusion [see the orange dotted curve in Fig. 2(c) in the main text]. A -polarized plane wave with of 532 nm and an electric field of 1 V/m is incident with an angle of 55∘ relative to the sample normal. The perfectly matched layers surrounding the volume are set to absorb all outgoing waves. The dielectric constant of the Ag objects is referenced from the literature [68]. The surface normal component of the electric field is sampled.
An Ag flat plate is placed beneath the tip apex with a gap distance . Here we define the origin of the gap distance ( = 0) as the tip height when the Ag-tip apex forms point contact with the Ag surface. As shown in Table S3, the molecular height is calculated to be 0.32–0.33 nm for D2/Ag(111). In the definition, taking into account the atomic radius of Ag, the position of the molecule corresponds to 0.2 nm above the Ag surface plane in the FEM model. Although the calculated depends on the tip height (Table S3), in the FEM model, we used the constant height for the sampling point at any gap distances. We confirmed that differences in the definition or the molecular position on the order of 0.1 nm do not result in qualitative differences in the field enhancement factor curves. To directly compare with the experimental TERS intensities, the electric field enhancement factor (Ref. [57]) is displayed in the plots of Fig. 2(c) in the main text. To the field enhancement of the nearest neighboring molecule, assuming the molecule is placed in the next adsorption site of Ag(111), = 0.29 nm is used [red triangle in the inset of Fig. 2(c) in the main text]. Thus, when the point at the surface just under the tip is defined as (, , ) = (0, 0, 0), the black and red curves shown in Fig. 2c of the main text represent the plots sampling at (, , ) = (0, 0, 0.20) nm and (0.29, 0, 0.20) nm, respectively.
As shown in Fig. 2, the gap distances with the FEM simulations pm to 850 pm well reproduced the Raman enhancement at the experimental tip height to 0 pm. Considering the Ag-atom radius, for the simulations (Fig. 4 in the main text) is comparable to nm. Therefore, pm corresponds to nm, which was used for the alignment of the experimental and theoretical plots displayed in Fig. 4(e) of the main text.
II II. Supplemental Results and Discussion
II.1 A. Nuclear Spin Isomers
H2 and D2 have two nuclear spin isomers, ortho and para. para-H2 (nuclear spin quantum number = 0) is more stable than ortho-H2 ( = 1) by 14 meV [69]. para-H2 has only even rotational quantum numbers , whereas and ortho-H2 has only odd . Although ortho-H2 is the predominant species in the gas phase at room temperature (a temperature equilibrium composition of 75%), the molecular adsorption on metals enhances the ortho-to-para conversion, leading to the dominance of the para species on the surfaces [4]. In the TERS spectra, we can detect only -H2 (Table S4); as a reference, in the gas phase, ortho-H2 have a Raman-active rotational mode [] at 587 cm-1 [69, 56]. This implies that we detected only the physisorbed species after ortho-para conversion on the surface. In the case of D2, ortho ( = 0, 2) is more stable than para ( = 1) by 7.4 meV and more dominant (67%) in the gas phase at room temperature. Only -D2 was detected in our TERS spectra, indicating again that para-D2 was converted on the surface.
Although we introduced normal hydrogen gases (n-H2 and n-D2) into the chamber, we observed rotational/vibrational transitions only for para-H2 and ortho-D2 in the TERS spectra (Fig. 1 in the main text). The dominance of para-H2 on Ag(111) in the TERS spectra ( = 10 K) is consistent with a previous HREELS study of H2 on an Ag film (exposures of 0.5–10 L, = 10 K) [8], whereas some previous reports detect ortho-H2 on Ag(111). A HREELS study on H2/Ag(111) at 6 K [11] shows that rotational mode peaks of both para-H2 and ortho-H2 were detected under a higher H2 pressure ( Torr) while para-H2 was dominantly detected under a lower pressure ( Torr). Photostimulated desorption studies [14, 15] shows that at an exposure of 30 L and a temperature of 6 K, ortho-H2 and para-D2 are alive on Ag(111) on the order of hundred seconds and that the conversion of D2 is slower than H2.
To compare with the above-mentioned studies, we performed TERS measurements while exposing D2 to a clean Ag(111) surfaces, as shown in Fig. S2. The gas pressure was set to Torr, comparable with the EELS study [11]. The ortho-D2-derived peaks appeared at 16 L and then saturated at 40 L, while para-D2 was not detected. The weak intensity at low exposures may be due to molecular diffusion along the surface at low coverages and the weak contribution of neighboring molecules to the TERS signal [see Fig. 2(c) in the main text]. It should be noted that even if the gas pressure and exposure values are comparable, the molecular coverage varies depending on the chamber configuration. Nevertheless, we confirmed that ortho-H2 and para-D2 were not detected at any pressures and exposures we used (0.8–2 Torr; 102–103 L). The HREELS observation of the metastable isomer at 6 K only at high exposures [11] may be attributed to be hydrogen multilayers, where the second layer does not interact with the Ag surface. For our LT-TERS measurement, in contrast, only a monolayer would be formed at the higher temperature, 10 K. Another possibility that the metastable isomer were not detected in TERS is that the LSPR in the gap is accelerating the para-to-ortho conversion of the molecule under the tip. Although far-field 532-nm irradiation did not affect the conversion rate [15], the strong electric field localized in the STM junction could interact with the nuclear spin of the adsorbate.
System | On Ag(111) | On Ag film | On Cu(001) | Gas phase | ||
Method | TERS (low res.) | TERS (high res.) | HREELSa | HREELSb | Ramanc | Ramand |
H2 | ||||||
rot | 351 | 348 | 395 | 363 | 354.5 | 354.365 |
vib | 4121 | 4134 | 4178 | 4178 | - | 4161.200 |
rot+vib | 4443 | - | 4533 | 4517 | - | 4497.848 |
D2 | ||||||
rot | 169 | 172 | - | 177 | 178.8 | 179.108 |
vib | 2967 | 2968 | - | 3000 | 2996.1 | 2993.6 |
rot+vib | 3133 | - | - | 3170 | 3166.3 | - |
II.2 B. Excitation Wavelength Dependence
To verify the contribution of LSPR to the TERS intensity, we examined the excitation wavelength () dependence. Figure S3(a) shows a STML spectrum recorded over a clean Ag(111) surface with an Ag tip. The LSPR results in a broad spectral feature in the visible range with a maximum at 560 nm. After recording the STML, we exposed the surface to the D2 gas, conducting TERS measurements with = 532 and 633 nm [Figs. S3(b) and S3(c), respectively]. The peak intensity of each mode significantly changes with , whereas the Raman shift value is unchanged. The vibrational intensity is weakened with = 633 nm, indicating the corresponding Raman scattering is not efficiently enhanced by the LSPR because the peak position ( = 779 nm) is located in the weak resonance range of the LSPR [Fig. S3(a)]. In contrast, both the rotational and vibrational peaks become intense with = 532 nm because they are more resonant with the LSPR.
II.3 C. Tip-Height Dependence of Peak Intensity and Width
To obtain the waterfall plots shown in Fig. 3 of the main text, we conducted tip approach-and-retraction procedures while recording the TERS spectra [54, 41]. Figure S10(a) is the full dataset for the H2 “vib” mode shown in Fig. 3(b) of the main text. During recording TERS spectra, the tip height was lowered stepwise and then the tip was retracted from the surface. The symmetrical waterfall plot with respect to the closest tip height indicates the non-destructive measurement in the tip approach-and-retraction process. Figure S10(b) shows the horizontal line profiles of the waterfall plot for the H2 “vib” mode at the points indicated by the arrows in the same color. The three spectra indicates the red-shift and broadening of the H2 “vib” peak as the gap distance decreases.
During recording the waterfall TERS plot, we occasionally observed a sudden intensity change caused presumably by the modification of the tip apex structure due to the strong tip–surface interaction at the small gap. Figure S10(c) shows an example, where a sudden drop of the TERS intensity was observed. The modification of the plasmon background intensity implies that the atomic-scale structure was changed by the strong attractive force exerted when the tip was brought close to the surface. At the same time as the plasmon feature, the D2 “rot” peak was also dropped, while the frequency of the “rot” mode is unchanged. This observation is in good agreement with the enhancement mechanism described in the main text (Fig. 2); the picocavity-derived electric field plays a crucial role in the TERS intensity of physisorbed hydrogen molecules.
To determine the experimental peak position shift as a function of the tip height in Fig. 4(e) of the main text, the line profile of the waterfall plots at each tip height [Fig. 3 in the main text; see also Fig. S10(b)] was fitted by a Gaussian peak with a constant background. The top and bottom panels of Fig. S11 shows the peak position and width, respectively, as a function of the tip height for the rotational (a) and vibrational (b) modes. The error bars indicate the standard deviation of the fitting. As the tip approaches, the peak intensity increases, reducing the error. The point at which the error becomes below 2 cm-1 by the tip approach was used as the reference point [the arrows in the top panel of Fig. S11(a) and (b)] and the peak position at this reference point was set to the origin of the peak shift (0 cm-1 of the top panel). For the H2 and D2 “vib” modes [top panel of Fig. S11(b)], the plots after the reference points were used as the experimental plots shown in Fig. 4(e) of the main text. In Fig. 4(e) of the main text, the experimental tip height pm is aligned to the tip–sample distance nm for the calculations. Although the accurate determination of the experimental tip–sample gap distance is difficult, this alignment is reasonable; as described in Sect. I-E, pm is comparable with pm and nm.
As shown in Fig. S11(a), the frequency and width of the H2 and D2 “rot” peaks are almost unchanged by the tip. In contrast, the bottom panel Fig. S11(b) shows the strong peak broadening for the H2 “vib” mode [see also Fig. S10(b)] whereas the peak width for D2 “vib” is almost unchanged. A possible origin of the peak broadening is the anharmonicity, which can modify the lifetime of molecular vibrations [71]. It is also entirely possible that if neighboring molecules influence the experimental signal [see Fig. 2(c) in the main text], they would lead to the high-frequency shoulder of the vibration transition for H2, as these molecules would still lie on the surface [Fig. 3(c)] and would therefore present a higher than the ones on the tip, which instead provide the most enhanced signal; however, this effect may not be pronounced because the coverage of is less than that of D2 (see Sect. I-C). Another possible origin if the movement of the molecule in the junction; because the double-well potential barrier is very small [see Fig. 4(d) in the main text], the H2 molecule near the tip apex can move to a metastable site near the surface or adjacent adsorption sites on the surface [18] by a small voltage bias or plasmonic field applied in the STM junction. This movement instantaneously occurring during the acquisition time of the Raman spectrum can also modify the peak shape of the vibrational mode. We expect that the coexistence of these factors leads to the characteristic peak width depending on the tip height.